HUBBLE SPACE TELESCOPE DAILY REPORT #926 PERIOD COVERED: 0800 EST Daily Status Report as of 193/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/BL 3618 (Execution Processes for the Outer Planet UV Emissions: Future - Cycle C) The Faint Object Spectrograph (Blue Side) was used to observe the planet Saturn to learn more about the processes responsible for the "electroglow" emission in the planet's upper atmosphere. Blind pointing was used to acquire the target and course track was used for guiding, since diffuse emission was being observed and the proposes want to avoid loss of lock. The observations were completed in fine lock . Some of the observations could not be displayed due to power problems at Goddard Space Flight Center (See Note 2.4). The observations will be analyzed later from playback tapes. 1.2 Completed PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) Rapid Internal Monitoring observations were made with both the Wide Field Camera and the Planetary Camera using filter F555W to look for new forms of contamination (i.e. daises and measles), and the possible return of quantum efficiency hysteresis. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.3 Completed FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3: cycle 3 High) The Faint Object Spectrograph was used to make one observation of quasar 2215-0347 to look for absorption lines at low redshift, and also emission lines in the QSO's continuum. The observation was executed as scheduled, and no problems were reported. The spectral observations was reported to look good. 1.4 Completed FOS RD 4396 (QSO Absorption System Snapshot Survey - Part 1 of 3: Cycle 3 High) The Faint Object Spectrograph (Red Side) was used to make two sets of spectrographic observation of quasar 0100+0205 and 0117+2118 to look for absorption line systems at low red shift, emission lines, and to study the QSO's continuum. The first set of observations could not be monitored due to a power outage at GSFC. (See Note 2.4) Nevertheless, the observations were later analyzed and reported to look fine. The second set of observations were also reported to look fine. 1.5 Completed Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of two observations were received. Both observations were reported to be saturated. 1.6 Completed Three Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make two standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. The first observation was partially saturated, and the second was completely saturated. 1.7 Completed FOC/96 4113 (Auroral Imaging of Jupiter with the FOC Cycle 2/ February-March 1993) The Faint Object Camera was used to make two images of the auroral emissions on the planet Jupiter at 1600 A0. Both north and south observations were contained within the sequence. The OSS verified that the correct filter was in place, but a rejected TDRSS contact prevented the OSS from uplinking a pointing correction until after the first observation had occurred. The first observation did not point at Jupiter. After the maneuver was completed, the subsequent observations were received and displayed. They were reported to be well centered, and to be well exposed. 1.8 Completed Two Set of FOC/96 4821 (MAGELLANIC Cloud Planetary Nebulae: Cycle 3 Med.) The Faint Object Camera (f/96) was used to make one observation of the planetary nebulae LMC-N66 which is located in the Large Magellanic Cloud to determine its mass and age. The observations were executed as scheduled, and no problems were reported. 1.9 Completed Two Set of WFC 4849 (GTO High Latitude Parallel Program: Cycle 3 High) The Wide Field Camera was used to make high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as planned, and no problems were reported. 1.10 Completed GHRS/2 4876 (Weak Interstellar Absorption Toward Zeta Orp - Augmentation Cycle 3 High) The High Resolution Spectrograph was used to make observations of the star Zeta Oph (Visual Mag. of 2.56). This star has interstellar absorption lines, and was observed using the Echelle Mode spectrograph to obtain spectra fo the weak interstellar line of Silicon at 3235 A0. The OSS monitored the observations and assessed the data quality. No problems were reported. 1.11 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.12 Completed GHRS/2 4272 (Supersonic Chromoshperic Winds - Cycle 3 Med) The High Resolution Spectrograph was used to make observations of the star Alpha Aquari (Visual Mag. of 2.90) to learn more about the chromosphereic winds by looking at the fluxes, profiles, and shifts of several spectral lines which are produced in the transition region where the winds are accelerated. The observations were completed as scheduled, and no problems were reported. 1.13 Completed Two Sets of FOS/BL 3545 (The Evolution of the UV Spectra in Early Type Galaxies Out to Z = 0.7) The Faint Object Spectrograph was used to make two sets of observations of the weak radio galaxy 53W022. These observations were executed in parallel with other science observations. Both sets of observations were received and displayed. No anomalies were noted. 1.14 Completed FOC/96 4969 (Search for QSOs Suitable for Subsequent Observations of He II) The Faint Object Camera was used to make ultraviolet images of the quasi-stellar object 1333+2821 to measure the levels of its ultraviolet flux to see if this object is suitable for subsequent observation with the FOS and GHRS. They are looking for absorption features of Helium II at 304 A0. There was a failed FHST update prior to the observation, but the guide star acquisition did achieve fine lock on a star. (See Note 2.2) The subsequent observations showed no target. 1.15 Completed FOC/96 4804 (UV Imaging of Nearby Galaxies with the FOC/96) The Faint Object Camera was used to complete one snapshot ultraviolet observation of the nearby galaxy UGC-11012 The exposures lasted for ten minutes and used filter F220W. The observations was reported to look fine. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 7 Successful : 7 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 16 Successful : 14 Failures at 192/1933Z and 192/1936Z, FHST 1 & 2 sensed errors, errors were not corrected. Subsequent acquisitions were successful, and following maps showed small errors. 2.3 Significant HSTARs: None 2.4 Operations Note: A major power outage occurred on Friday 190/1511Z in the Building 3/14 area when electrical power was switched for commercial power to diesel power. The power outage was building wide. The equipment automatically switched over to the backup UPS System, and ran on batteries for about 20 minutes, until the batteries ran out. The power outage lasted for about 1 hour 19 minutes. Two periods of 26 minutes 21 seconds, and 6 minutes 5 seconds of real-time telemetry were lost and are unrecoverable. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Set of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.2 Uplink Updated Ephemeris Tables , Per MOSES 2871 3.3 Complete FOS/BL 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle 3 High) 3.4 Complete FOS/BL 4661 (Ultraviolet Eclipsing Mapping of the Accretion Flow in the Eclipsing Int.) 3.5 Complete Two Set of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.6 Complete Two Set of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.7 Complete WFC PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-